# Phase diagram and universal scaling regimes of two-dimensional exciton–polariton Bose–Einstein condensates

### v1

@@ IV.3 Steady-state condensate density
Results are shown as a function of the blueshift at threshold on the upper panel of Fig. 2, revealing a sharp transition at $\mu_{\rm th}\approx 480{\rm\mu eV}$. This abrupt drop of density is related to the transition from a vortex-free condensate with very few space-time vortices to a defect-dominated phase (blue-shaded area). A similar transition was reported in Ref. [35] as a result of modulational instability [78, 79, 42, 26], whereas in the present case it rather seems to be related to the competition between the finite system size and the typical vortex separation $L_{v}$ scale [39], establishing a transition from a phase turbulent to a defect turbulent regime [70].
Figure 2: Condensate density averaged over space and $128$ realizations of the noise, normalized by $n_{0}$ (upper panel) and effective KPZ nonlinearity defined from Eq. (3) (lower panel) as a function of the blueshift at threshold $\mu_{\rm th}$. Light, medium, and dark blue circles indicate points of the phase diagram corresponding to the EW, KPZ, and vortex regime respectively.
@@ IV.4 First-order correlation $g^{(1)}(r,t)$
Each of the regimes presented in Sec. III deeply imprint the phase dynamics of the EP condensate, and therefore its coherence properties. They can be distinguished experimentally in the spatiotemporal decay of the condensate first-order correlation function, defined as
@@
[…]
@@
$-2\log\left[|g^{(1)}_{\psi}(\Delta r,\Delta t)|\right]\simeq C_{\theta\theta}(\Delta r,\Delta t)$ (11)
@@
Figure 3: Equal-space (upper panel) and equal-time (lower panel) first-order correlation function corresponding to the EW, KPZ, and vortex points of Fig. 2. A logarithmic fit is plotted in dotted line for the former, and guidelines are shown in dash-dotted and dashed line for the two latters. The red horizontal dashed line corresponds to the expected saturation value of $g^{(1)}(0,\Delta t)$ in the absence of density correlations and for $\text{Var}\left[\theta(r_{0},\Delta t+t_{0})-\theta(r_{0},t_{0})\right]=2\pi$. The local exponents, computed acording to Eqs. (14a)–(14b), are shown for the KPZ point in the insets, where the grey shade extends, as the dash-dotted line, over $\left[35{\rm ps},350{\rm ps}\right]$ and $\left[15{\rm\mu m},50{\rm\mu m}\right]$ respectively.
@@
$\mathcal{G}^{(1)}(\Delta r,\Delta t)\sim\left\{\begin{array}[]{l l}\Delta t^{2\beta_{\rm KPZ}},&\Delta r=0\\ \Delta r^{2\alpha_{\rm KPZ}},&\Delta t=0\end{array}\right.\,,$ (12)
@@
as suggested in early studies of the CGLE [71]. As already mentioned in Sec. I, this behavior has been confirmed in experimental realizations of 1d polariton condensates [24], as well as in numerical simulations of their 2d analog [8, 76].
@@
For weak but non-vanishing nonlinearity, EW correlations are expected to persist for extended scales before KPZ features arise. While stretched exponential decays of $g^{(1)}$ are expected in 1d in the EW phase as in the KPZ phase [9], the situation changes in 2d as a result of the vanishing of $\alpha_{\rm EW}$ and $\beta_{\rm EW}$ (Sec. III.2). This yields instead logarithmic growth as predicted from the solution of the EW equation [44]
@@
$\mathcal{G}^{(1)}(\Delta r,\Delta t)\sim\left\{\begin{array}[]{l l}\dfrac{D}{2\pi\nu}\ln(\Delta t\Lambda^{2}),&\Delta r=0\\ \dfrac{D}{\pi\nu}\ln(\Delta r\Lambda),&\Delta t=0\end{array}\right.\,,$ (13)
@@
with $\Lambda$ a UV cutoff. This implies an algebraic decay of $g^{(1)}$ in space as $g^{(1)}(\Delta r,0)\sim\Delta r^{-a_{s}}$ with exponent $a_{s}=D/2\pi\nu$, and in time as $g^{(1)}(0,\Delta t)\sim\Delta t^{-a_{t}}$ with exponent $a_{t}=D/4\pi\nu$. Such a behavior was predicted within Bogoliubov theory [81, 82] and observed in experiments [83] but however never related to EW universal features. This identification, motivated by the mapping (3) as well as results shown on Fig. 2, appears crucial in guiding the observation of KPZ universality in 2d polariton condensates. Moreover, this result provides a quantitative explanation for the exponent ratio estimated numerically as $a_{s}\simeq 2a_{t}$ [34], as well as for the recent experimental observation [84] that $a_{s}\propto 1/n_{0}$ from which it was suggested $a_{s}\propto\frac{\sigma/n_{0}}{\hbar/m}$ in analogy with equilibrium condensates. Indeed, (13) captures this behavior through $a_{s}\propto\frac{D}{\nu}$, where the nonequilibrium nature of the polariton condensate appears explicitly in the effective noise (6) and viscosity (4) of the phase. The general correspondence between Bogoliubov prediction for the large space and time behavior of the phase correlation and EW properties is detailed in Appendix A.
@@
Figure 4: Equal-space (upper panel) and equal-time (lower panel) second-order correlation function.
@@
The first-order correlation function is computed in the steady state of the condensate according to Eq. (8), and averaged over 1024 realizations of the noise. The results obtained in the three regimes are shown on Fig. 3. In the weak nonlinearity regime (light blue), algebraic decays consistent with EW behavior (13) are observed, and the corresponding fits are shown as dotted line, where we find $a_{s}/a_{t}=1.8\pm 0.1$. These fits are performed in a restricted spatiotemporal window (grey-shaded area). Its lower bound corresponds to the time where density correlations, displayed in Fig. 4, are negligible ($\mathcal{G}^{(1)}_{n}$ is flat) such that the first-order correlation function is fully determined by the phase. Its upper bound corresponds to the time when $g^{(1)}(0,\Delta t)$ develops an exponential decay induced by the finite system size (discussed in Sec. IV.5).
@@
At intermediate values of $\mu_{\rm th}$ (medium blue), a clear departure from EW behavior is observed. This departure is induced by phase correlations, as shown on Fig. 4, where the decay of density correlations perfectly matches the one at the EW point, indicating that approximations leading to the effective phase dynamics (3) are valid. Reservoir effects are visible at times $\Delta t\approx 10$ to $30$ps and naturally arise when interactions are increased or for high values of the pumping rate [26, 76]. Proceeding as in Ref. [76], we compute the local exponents according to
@@ IV.5 Finite-size effects
The spatio-temporal windows where EW or KPZ critical exponents can be measured are very small, which hinders the observability of these regimes. However, interestingly, these exponents also imprint the exponential decay superseding the EW-KPZ regimes, through finite-size effects. We thus now investigate this exponential decay, which can be observed on Fig. 3 in the equal-space first-order correlation function at the latest times. For this, we compute the condensate first-order correlation function averaged over 256 realizations of the noise at the EW point of Fig. 2 increasing the system size $N\times N$ from $N=8$ to $N=256$ for temporal correlations and up to $N=1024$ for spatial correlations. Results are shown on Fig. 5, where one observes that the condensate coherence preserves a quasi long-range order down to millimetre lengths. As a function of time, the condensate exhibits quasi long-range order up to a crossover time, that depends on the system size, from which the decay becomes exponential. This result can be summarized in the following scaling form for the equal-space phase-phase correlator
@@
which was first formulated in the context of kinetically roughening interfaces [85], then suggested in the context of the CGLE [71], and very recently again in the large time decoherence of 1d driven-dissipative condensates [86]. A derivation of Eq. (15) is presented within Bogoliubov theory in both one and two dimensions in Appendix A. This equation represents the change in behavior of correlations between their thermodynamics limit and their saturation to Brownian motion, when the growing interface (here the phase) looses the memory of its initial condition [85].
@@
Figure 5: First-order correlation function at the EW point of Fig. 2 for a system of $N=2^{i}$ linear sites, at (upper panel) equal-space with $i=\{3,\dots,8\}$, and (lower panel) equal-time with $i=\{3,\dots,10\}$. The colorscale is growing with the system size from light green to dark blue. Algebraic decays are represented by the fitted orange dashed line. A grey dotted line indicates the large time exponential decay. Inset: coherence time $\uptau_{c}$ fitted from the large-time exponential decay of $g^{(1)}(0,\Delta t)$ as a function of the system size $L$.
@@
Figure 6: First-order correlation function at the KPZ point of Fig. 2 for a system of $N=32$ to 64 linear sites, at (upper panel) equal-space, and (lower panel) equal-time. The colorscale is growing with the system size from light green to dark blue. Algebraic decays are represented by the dash-dotted line. A grey dotted line indicates the large time exponential decay. Inset: coherence time $\uptau_{c}$ fitted from the large-time exponential decay of $g^{(1)}(0,\Delta t)$ as a function of the system size $L$.
When the nonlinearity $\lambda$ is large enough to let KPZ features arise, we expect $\tau_{c}\sim L^{0.83}$. Proceeding in the same manner as above, we compute the condensate first-order correlation function averaged over 2048 realizations at the KPZ point of Fig. 2, varying N. As we have discussed in Sec. IV.3, the position of the transition in Fig. 2 depends on the size of the system, causing the phase turbulence regime where KPZ features are observed to disappear in favour of defect turbulence if the system size is too large. For this reason, N is varied from 32 to 64 only. Results are shown on Fig. 6, where the spatial decoherence (lower panel) visibly converges to the KPZ scaling as the system size increases, as observed in Ref. [8]. The temporal decoherence (upper panel) shows that KPZ features are dramatically affected when the system gets smaller, the scaling window becoming hardly discernible. However, its signature remains accessible from the coherence time according to Eq. (15), whose dependence on the system size is close to linear (Inset of Fig. 6) and clearly differs from the parabola obtained on Fig. 5. Fitting the numerical values of the coherence time with a power law ansatz, we find $\uptau_{c}\sim L^{0.97\pm 0.18}$, which closely encloses the prediction from KPZ exponents and discard the one from EW exponents.
@@ IV.6 Momentum distribution $n_{\hat{\bm{k}}}$
where $\psi(\hat{\bm{k}},t_{0})$ is the Fourier transform of the condensate wave function. The results are displayed on Fig. 7 for the three points shown on Fig. 2. Let us first comment on the behavior of the momentum distribution for the large values of $k$. In the adiabatic approximation of the reservoir, Bogoliubov theory predicts in the continuous limit that the momentum distribution is determined by the microscopic properties of the condensate according to $n_{\hat{\bm{k}}}\propto\sigma/\gamma_{2}k^{2}$ (see Appendix A). We therefore expect the three curves to share the same decay at large $k$. This is indeed observed, and the momentum distribution decays as $n_{\hat{\bm{k}}}\sim k^{-1.893\pm 0.004}$, represented on Fig. 7 by a dashed guideline, in good agreement with the theoretical prediction.
@@
At small momenta, we expect $n_{\hat{\bm{k}}}$ to depend on the condensate macroscopic properties, so that it should behave differently in the three regimes under consideration. Within the conditions of validity of Bogoliubov theory, in particular for small density fluctuations, we find $n_{\hat{\bm{k}}}\propto D/\nu k^{2\beta z+2}$. This leads to the prediction $n_{\hat{\bm{k}}}\propto D/\nu k^{2}$ at the EW point. On Fig. 7, we rather observe $n_{\hat{\bm{k}}}\sim k^{-1.893\pm 0.004}$ as in the large $k$ regime, which is in good agreement with the theoretical expectation. At the KPZ point, nonlinearity comes into play and the scaling is expected to change to $n_{\hat{\bm{k}}}\sim k^{-2.77}$. The observed behavior is fairly close to this prediction, as we obtain $n_{\hat{\bm{k}}}\sim k^{-2.93\pm 0.03}$ as shown by the dotted guideline.

### v3

@@ IV.3 Steady-state condensate density
Results are shown as a function of the blueshift at threshold on the upper panel of Fig. 2, revealing a sharp transition at $\mu_{\rm th}\approx 480{\rm\mu eV}$. This abrupt drop of density is related to the transition from a vortex-free condensate with very few space-time vortices to a defect-dominated phase (blue-shaded area). A similar transition was reported in Ref. [35] as a result of modulational instability [88, 89, 43, 26], whereas in the present case it rather seems to be related to the competition between the finite system size and the typical vortex separation scale $L_{v}$ [40], establishing a transition from a phase turbulent to a defect turbulent regime [71].
Figure 2: Condensate density averaged over space and 128128 realizations of the noise, normalized by $n_{0}$ (upper panel) and effective KPZ nonlinearity defined from Eq. (3) (lower panel) as a function of the blueshift at threshold $\mu_{\rm th}$. Light, medium, and dark blue circles indicate points of the phase diagram corresponding to the EW, KPZ, and vortex regime respectively.
@@ IV.4 First-order correlation $g^{(1)}(r,t)$
Each of the regimes presented in Sec. III deeply imprints the phase dynamics of the EP condensate, and therefore its coherence properties. They can be distinguished experimentally in the spatiotemporal decay of the condensate first-order correlation function, defined as
@@
where the average $\langle\cdot\rangle$ is performed over realizations of the stochastic dynamics (1) in the steady state, as well as over the circle of radius $\Delta r=|\Delta\bm{r}|$ exploiting isotropy. This correlation function encompasses contributions from both the density and the phase, and approximately reads
@@
$2\log\left[|g^{(1)}_{\psi}(\Delta r,\Delta t)|\right]\simeq C_{\theta\theta}(\Delta r,\Delta t)$ (11)
@@
Figure 3: Equal-space (upper panel) and equal-time (lower panel) first-order correlation function corresponding to the EW, KPZ, and vortex points of Fig. 2. A logarithmic fit is plotted in dotted line for the former, and guidelines are shown in dash-dotted and dashed line for the two latters. The red horizontal dashed line corresponds to the expected saturation value of $g^{(1)}(0,\Delta t)$ in the absence of density correlations and for $\text{Var}\left[\theta(r_{0},\Delta t+t_{0})-\theta(r_{0},t_{0})\right]=2\pi$. The local exponents, computed according to Eqs. (14a)–(14b), are shown for the KPZ point in the insets, where the grey shade extends, as the dash-dotted line, over $\left[35{\rm ps},350{\rm ps}\right]$ and $\left[15{\rm\mu m},50{\rm\mu m}\right]$ respectively.
@@
$\mathcal{G}^{(1)}(\Delta r,\Delta t)\sim\left\{\begin{array}[]{l l}\Delta t^{2\beta_{\rm KPZ}},\qquad&\Delta r=0\\ \Delta r^{2\alpha_{\rm KPZ}},&\Delta t=0\end{array}\right$.\,, (12)
@@
as suggested in early studies of the CGLE [72]. As already mentioned in Sec. I, this behavior has been confirmed in experimental realizations of 1d polariton condensates [24], as well as in numerical simulations of their 2d analog under idealized conditions [8, 38].
@@
For weak but non-vanishing nonlinearity, although the system is out-of-equilibrium, EW correlations are expected to persist for extended scales before KPZ features arise. While stretched exponential decays of $g^{(1)}$ are expected in 1d in the EW phase as in the KPZ phase [9], the situation changes in 2d as a result of the vanishing of $\alpha_{\rm EW}$ and $\beta_{\rm EW}$ (Sec. III.2). This yields instead logarithmic growth as predicted from the solution of the EW equation [45]
@@
$\mathcal{G}^{(1)}(\Delta r,\Delta t)\sim\left\{\begin{array}[]{l l}\dfrac{D}{2\pi\nu}\ln(\Delta t\Lambda^{2}),\qquad&\Delta r=0\\ \dfrac{D}{\pi\nu}\ln(\Delta r\Lambda),&\Delta t=0\end{array}\right$.\,, (13)
@@
with $\Lambda$ a UV cutoff. This implies an algebraic decay of $g^{(1)}$ in space as $g^{(1)}(\Delta r,0)\sim\Delta r^{-a_{s}}$ with exponent $a_{s}=D/2\pi\nu$, and in time as $g^{(1)}(0,\Delta t)\sim\Delta t^{-a_{t}}$ with exponent $a_{t}=D/4\pi\nu$. Such a behavior was predicted within Bogoliubov theory [91, 92] and observed in experiments [93] but however never related to EW universal features. We highlight that the EW power-law exponents are different from the 2d closed case, where the BKT ordered phase displays an exponent ratio $a_{s}/a_{t}=1$. Values $a_{s}/a_{t}\simeq 1$ were reported in experiments on quasi-equilibrium polariton condensates [29]. The identification of the EW regime, motivated by the mapping (3) as well as results shown on Fig. 2, appears crucial in guiding the observation of KPZ universality in 2d polariton condensates. Moreover, this result provides a quantitative explanation for the exponent ratio estimated numerically as $a_{s}\simeq 2a_{t}$ in Ref. [34], as well as for the recent experimental observation [79] that $a_{s}\propto 1/n_{0}$.
Let us note that in Ref. [79], it was suggested that $a_{s}\propto 1/n_{0}\lambda_{T}^{2}$ in analogy with equilibrium condensates with $\lambda_{T}=\hbar/\sqrt{2\pi mk_{B}T}$ the thermal de-Broglie wavelength. This expression is however ambiguous as a temperature $T$ is not defined out of thermal equilibrium. An effective temperature can however be rigorously defined, comparing the tails of the polariton condensate momentum distribution with that of a thermal closed Bose gas [91], where in our notations $k_{B}T_{\rm eff}=\frac{2D}{\nu}\frac{\hbar^{2}n_{0}}{2m}$. The spatial EW exponent can in turn be rewritten $a_{s}=1/n_{0}\lambda_{T_{\rm eff}}^{2}$, making the non-equilibrium nature of the condensate perfectly explicit since the effective temperature depends on drive and dissipation, and the noise is not constrained by any fluctuation-dissipation type of relation. Note that to take into account actual thermal effects, an additional phonon bath can be included and it is found to contribute to the noise strength $\sigma$ [94, 95]. The general correspondence between Bogoliubov prediction for the large space and time behavior of the phase correlation and EW properties is detailed in Appendix A.
@@
Figure 4: Equal-space (upper panel) and equal-time (lower panel) density correlation.
@@
The first-order correlation function is computed in the steady state of the condensate according to Eq. (8), and averaged over 1024 realizations of the noise. The results obtained in the three regimes are shown on Fig. 3. In the weak nonlinearity regime (light blue), algebraic decays consistent with EW behavior (13) are observed, and the corresponding fits are shown as dotted line, where we find $a_{s}/a_{t}=1.8\pm 0.1$. These fits are performed in a restricted spatiotemporal window (grey-shaded area). Its lower bound corresponds to the time where density correlations (second-order correlation function), displayed in Fig. 4, are negligible ($\mathcal{G}^{(1)}_{n}$ is flat) such that the first-order correlation function is fully determined by the phase. Its upper bound corresponds to the time when $g^{(1)}(0,\Delta t)$ develops an exponential decay induced by the finite system size (discussed in Sec. IV.5).
@@
At intermediate values of $\mu_{\rm th}$ (medium blue), a clear departure from EW behavior is observed. This departure is induced by phase correlations, as shown on Fig. 4, where the decay of density correlations perfectly matches the one at the EW point, indicating that approximations leading to the effective phase dynamics (3) are valid. Reservoir effects are visible at times $\Delta t\approx 10$ to 3030ps and naturally arise when interactions are increased or for high values of the pumping rate [26, 38]. Proceeding as in Ref. [38], we compute the local exponents according to
@@ IV.5 Finite-size effects
In addition to the scaling laws expected in the decay of the coherence, we show below that KPZ or EW critical exponents also imprint the exponential decay superseding the EW-KPZ regimes, through finite-size effects. We thus now investigate this exponential decay, which can be observed on Fig. 3 in the equal-space first-order correlation function at the latest times. For this, we compute the condensate first-order correlation function averaged over 256 realizations of the noise at the EW point of Fig. 2 increasing the system size $N\times N$ from $N=8$ to $N=256$ for temporal correlations and up to $N=1024$ for spatial correlations. Results are shown on Fig. 5, where one observes that the condensate coherence preserves a quasi long-range order down to millimeter lengths. As a function of time, the condensate exhibits quasi long-range order up to a crossover time, that depends on the system size, from which the decay becomes exponential. This result can be summarized in the following scaling form for the equal-space phase-phase correlator
@@
which was first formulated in the context of kinetically roughening interfaces [96], then suggested in the context of the CGLE [72], and very recently again in the large time decoherence of 1d driven-dissipative condensates [97]. A derivation of Eq. (15) is presented within Bogoliubov theory in both one and two dimensions in Appendix A. This equation represents the change in behavior of correlations between their thermodynamics limit and their saturation for finite size, after the growing interface (here the phase) has lost the memory of its initial condition [96].
@@
Figure 5: First-order correlation function at the EW point of Fig. 2 for a system of $N=2^{i}$ linear sites, at (upper panel) equal-space with $i=\{3,\dots,8\}$, and (lower panel) equal-time with $i=\{3,\dots,10\}$. The colorscale is changing with the increasing system size from light green to dark blue. Algebraic decays are represented by the fitted orange dashed line. A grey dotted line indicates the large time exponential decay. Inset: coherence time $\uptau_{c}$ fitted from the large-time exponential decay of $g^{(1)}(0,\Delta t)$ as a function of the system size $L$.
@@
Figure 6: First-order correlation function at the KPZ point of Fig. 2 for a system of $N=32$ to 64 linear sites, at (upper panel) equal-space, and (lower panel) equal-time. The colorscale is changing with increasing the system size from light green to dark blue. Stretched exponential decays are represented by the dash-dotted line. A grey dotted line indicates the large time exponential decay. Inset: coherence time $\uptau_{c}$ fitted from the large-time exponential decay of $g^{(1)}(0,\Delta t)$ as a function of the system size $L$.
When the nonlinearity $\lambda$ is large enough to let KPZ features arise, we expect $\tau_{c}\sim L^{0.83}$. Proceeding in the same manner as above, we compute the condensate first-order correlation function averaged over 2048 realizations at the KPZ point of Fig. 2, varying N. As we have discussed in Sec. IV.3, the position of the transition in Fig. 2 depends on the size of the system, causing the phase turbulence regime where KPZ features are observed to disappear in favor of defect turbulence if the system size is too large. For this reason, N is varied from 32 to 64 only, including more intermediate points to improve the fit of the linewidth $\uptau_{c}(L)$. Results are shown on Fig. 6, where the spatial decoherence (lower panel) visibly converges to the KPZ scaling as the system size increases, as observed in Ref. [8]. The temporal decoherence (upper panel) shows that KPZ features are dramatically affected when the system gets smaller, the scaling window becoming hardly discernible. However, its signature remains accessible from the coherence time according to Eq. (15), whose dependence on the system size is close to linear (Inset of Fig. 6) and clearly differs from the parabola obtained on Fig. 5. Fitting the numerical values of the coherence time with a power law ansatz, we find $\uptau_{c}\sim L^{0.97\pm 0.18}$, which closely encloses the prediction from KPZ exponents and discard the one from EW exponents.
@@ IV.6 Momentum distribution $n_{\hat{\bm{k}}}$
where $\psi(\hat{\bm{k}},t_{0})$ is the Fourier transform of the condensate wave function and $t_{0}$ is chosen in the stationary state. The results are displayed on Fig. 7 for the three points shown on Fig. 2. Let us first comment on the behavior of the momentum distribution for the large values of $k$. In the adiabatic approximation of the reservoir, Bogoliubov theory predicts in the continuous limit that the momentum distribution is determined by the microscopic properties of the condensate according to $n_{\hat{\bm{k}}}\propto\sigma/\gamma_{2}k^{2}$ (see Appendix A). We therefore expect the three curves to share the same decay at large $k$. This is indeed observed, and the momentum distribution is found to decay as $n_{\hat{\bm{k}}}\sim k^{-1.893\pm 0.004}$, represented on Fig. 7 by a dashed guideline, in good agreement with the theoretical prediction.
@@
At small momenta, we expect $n_{\hat{\bm{k}}}$ to depend on the condensate macroscopic properties, so that it should behave differently in the three regimes under consideration. Within the conditions of validity of Bogoliubov theory, in particular for small density fluctuations, we find $n_{\hat{\bm{k}}}\propto D/\nu k^{2\beta z+2}$. This leads to the prediction $n_{\hat{\bm{k}}}\propto D/\nu k^{2}$ at the EW point. On Fig. 7, we observe $n_{\hat{\bm{k}}}\sim k^{-1.893\pm 0.004}$ as in the large $k$ regime, which is in good agreement with the theoretical expectation. At the KPZ point, nonlinearity comes into play and the scaling is expected to change to $n_{\hat{\bm{k}}}\sim k^{-2.77}$. The observed behavior is fairly close to this prediction, as we obtain $n_{\hat{\bm{k}}}\sim k^{-2.93\pm 0.03}$ as shown by the dotted guideline.